Home mework 2-1 lower Pot Falling Past a Window く) 2017 Constants I Periodic Table Express your answer in terms of some or all of the variables hb. Lw. t, h, and 9 As you look out of your dorm window, a flower pot suddenly falls past. The pot is visible for a time t and the vertical length of your window is Lw. Take down to be the positive direction, so that downward velocities are positive and the acceleration due to gravity is the positive quantity g View Available Hint(s) Assume that the flower pot was dropped by someone on the floor above you (rather tharn Viground ^ Zoorm launcher (1).exe IMG 2093 (2) jpg IMG 2094 (2)jpg
Answer #1

If we take the bottom of your window as the initial position and ground as the final position and use the kinematic equation, v^2 - u^2 =2as, we have, initial velocity, u = v_b, final velocity,v = v_ground, acceleration, a = g and displacent, s = h_b. Thus,

(v_ground)^2 - (v_b)^2 = 2*g*h_b

Or,    (v_ground)^2 = (v_b)^2 + 2*g*h_b

Thus, v_ground = \sqrt[(v_b)^2 + 2*g*h_b]